OpenStudy (anonymous):

What is the vertex of the following equation. I do not really understand how to do it, so if someone can show me a example of how to do it then I think I could do it. y = x^2 − 2x

OpenStudy (anonymous):

@rileychampion365

OpenStudy (anonymous):

If the equation is y = 3(x + 4)2 - 6, the value of h is -4, and k is -6. To convert a quadratic from y = ax2 + bx + c form to vertex form, y = a(x - h)2+ k, you use the process of completing the square. Let's see an example. Convert y = 2x2 - 4x + 5 into vertex form, and state the vertex.

OpenStudy (anonymous):

\[h= \frac{ -b }{ 2a }\]

OpenStudy (anonymous):

ok thanks. so for that equation I have I would have \[h=\frac{ -2}{ 2(2)}\]

OpenStudy (anonymous):

Mhmmm...

OpenStudy (anonymous):

How would I get the two points. h =4 what would my y axis be?

OpenStudy (anonymous):

To convert a quadratic from y = ax2 + bx + c form to vertex form, y = a(x - h)2+ k, you use the process of completing the square.

OpenStudy (anonymous):

I haven't learned that yet in algebra. So I would do 2(x-4)+0?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

wait so y would equal 2x-8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so then my point on a coordinate plane would be (2,-8) ?